OpenStudy (anonymous):

f(x) = 2xsqrtx+4/x+1, find g(-3)

OpenStudy (anonymous):

If g was not defined, g(-3) is an undefined value.

OpenStudy (anonymous):

\[f(x)=\frac{2x\sqrt{x+4}}{x+1}\] \[f(-3)=\frac{2\times (-3)\sqrt{-3+4}}{-3+1}\] is a start

OpenStudy (anonymous):

f(-3) = -6 sqrt7/-2?

OpenStudy (anonymous):

not quite inside the radical is \(-3+4=1\) not \(7\)

OpenStudy (anonymous):

and so the radical part becomes \(\sqrt{-3+4}=\sqrt 1=1\)

OpenStudy (anonymous):

I'm sorry. I had 3 + 4. I mistyped

OpenStudy (anonymous):

Thanks. You're a jewel

OpenStudy (anonymous):

yw (blush)